# Testing Low-Resource Language Support in LLMs Using Language Proficiency Exams: the Case of Luxembourgish

### v2

@@ C Language Exam Prompt
I will give you a language test for
@@
by [BLANK] and a list of possible ANSWERS
where every possible option is separated
with a comma. You have to choose the
option that best fits as replacement for
the given BLANK. As output, simply write
the chosen option, nothing else.
@@ D Evaluation Metric Prompts for LuxGen
    written without any grammatical or spelling
    mistakes. The more mistakes there are, the
    lower the score should be. A 0 means that
    the headline is full of mistakes and made
    up words. Also return 0 if the description
    is not written in Luxembourgish at all.
@@ E Coarse-grained Exam Results
A1 A2 B1 B2 C1 C2
Model Size Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC
Small Models $(\leq 15B)$
Qwen 2 0.5 38.5 26.9 23.1 19.2 34.6 11.5 19.2 23.1 32.0 15.4 23.1 23.1 25.9 15.2 19.2 30.8 19.2 28.6 15.4 20.0 8.0 34.6 11.5 19.2
Llama 3.2 1 0.0 0.0 3.8 3.8 11.5 3.8 0.0 0.0 4.0 0.0 0.0 0.0 11.1 3.0 0.0 3.8 30.8 14.3 0.0 0.0 4.0 7.7 3.8 0.0
Qwen 2 1.5 34.6 38.5 42.3 15.4 38.5 26.9 23.1 23.1 32.0 30.8 19.2 23.1 33.3 36.4 23.1 34.6 50.0 39.3 11.5 32.0 20.0 19.2 23.1 26.9
StableLM 2 1.6 0.0 0.0 0.0 0.0 19.2 23.1 0.0 15.4 0.0 0.0 0.0 0.0 25.9 24.2 11.5 3.8 19.2 7.1 0.0 12.0 24.0 26.9 0.0 11.5
Gemma 2 2 38.5 46.2 38.5 34.6 38.5 23.1 34.6 34.6 40.0 19.2 30.8 30.8 0.0 0.0 0.0 0.0 19.2 42.9 42.3 40.0 40.0 23.1 23.1 26.9
Llama 3.2 3 34.6 34.6 46.2 34.6 38.5 23.1 26.9 53.8 40.0 26.9 23.1 23.1 37.0 27.3 42.3 42.3 38.5 25.0 23.1 40.0 12.0 30.8 30.8 19.2
Phi 3 3.8 26.9 46.2 57.7 50.0 34.6 23.1 38.5 50.0 24.0 23.1 23.1 38.5 25.9 27.3 15.4 38.5 30.8 28.6 23.1 24.0 24.0 7.7 19.2 19.2
Phi 3.5 3.8 7.7 0.0 19.2 0.0 11.5 7.7 15.4 15.4 0.0 3.8 3.8 0.0 0.0 3.0 0.0 3.8 30.8 3.6 19.2 12.0 40.0 11.5 15.4 19.2
Qwen 2 7 42.3 26.9 69.2 53.8 23.1 38.5 38.5 46.2 36.0 23.1 15.4 61.5 29.6 33.3 38.5 42.3 38.5 32.1 38.5 36.0 36.0 42.3 15.4 38.5
Llama 2 7 7.7 0.0 7.7 15.4 0.0 0.0 19.2 23.1 8.0 0.0 23.1 3.8 7.4 6.1 7.7 34.6 26.9 17.9 23.1 28.0 24.0 15.4 23.1 19.2
WizardLM 2 7 7.7 15.4 38.5 34.6 30.8 26.9 30.8 34.6 24.0 34.6 19.2 30.8 7.4 18.2 26.9 26.9 38.5 32.1 26.9 44.0 16.0 34.6 15.4 26.9
Mistral 7 30.8 26.9 46.2 38.5 23.1 30.8 42.3 46.2 44.0 34.6 26.9 34.6 29.6 24.2 26.9 46.2 34.6 21.4 30.8 40.0 20.0 26.9 30.8 26.9
Aya-23 8 38.5 38.5 46.2 38.5 30.8 19.2 30.8 42.3 32.0 19.2 26.9 38.5 25.9 30.3 26.9 42.3 26.9 28.6 26.9 32.0 28.0 34.6 38.5 30.8
Llama 3.1 8 50.0 61.5 57.7 57.7 30.8 30.8 26.9 46.2 16.0 38.5 19.2 57.7 33.3 24.2 53.8 26.9 38.5 28.6 50.0 44.0 28.0 26.9 30.8 46.2
Llama 3 8 42.3 42.3 50.0 46.2 30.8 34.6 26.9 50.0 40.0 30.8 30.8 38.5 25.9 36.4 34.6 46.2 42.3 39.3 15.4 28.0 24.0 34.6 42.3 23.1
DeepSeek-R1 8 53.8 50.0 34.6 30.8 42.3 23.1 30.8 30.8 36.0 26.9 34.6 42.3 44.4 42.4 42.3 34.6 30.8 14.3 23.1 40.0 20.0 26.9 23.1 34.6
GLM 4 9 57.7 42.3 69.2 50.0 34.6 19.2 42.3 69.2 60.0 15.4 26.9 53.8 59.3 27.3 50.0 50.0 53.8 21.4 34.6 44.0 36.0 23.1 26.9 50.0
Gemma 2 9 76.9 50.0 61.5 46.2 61.5 38.5 30.8 57.7 52.0 26.9 34.6 46.2 0.0 0.0 0.0 0.0 38.5 17.9 46.2 48.0 36.0 23.1 42.3 30.8
Mistral-Nemo 12 61.5 50.0 53.8 46.2 34.6 26.9 38.5 46.2 44.0 15.4 23.1 38.5 33.3 24.2 38.5 46.2 30.8 25.0 26.9 32.0 20.0 26.9 26.9 11.5
StableLM 2 12 42.3 15.4 30.8 23.1 0.0 0.0 7.7 23.1 44.0 11.5 19.2 30.8 22.2 18.2 50.0 30.8 34.6 17.9 19.2 32.0 20.0 23.1 30.8 23.1
Llama 2 13 34.6 38.5 15.4 23.1 19.2 11.5 15.4 11.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 30.8 21.4 19.2 20.0 24.0 19.2 26.9 23.1
Phi 3 14 42.3 42.3 50.0 46.2 38.5 42.3 30.8 42.3 44.0 38.5 30.8 46.2 48.1 33.3 34.6 42.3 30.8 39.3 34.6 36.0 12.0 30.8 19.2 34.6
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 15.4 0.0 0.0 3.8 11.5 3.8 0.0 0.0 8.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 7.7 0.0 0.0 0.0 4.0 0.0 0.0 0.0
Gemma 2 27 42.3 23.1 3.8 15.4 15.4 11.5 0.0 0.0 0.0 7.7 0.0 7.7 3.7 3.0 0.0 0.0 7.7 3.6 0.0 0.0 4.0 3.8 3.8 0.0
Aya-23 35 15.4 7.7 0.0 7.7 7.7 0.0 0.0 7.7 4.0 0.0 0.0 0.0 7.4 3.0 0.0 3.8 15.4 0.0 0.0 4.0 4.0 0.0 0.0 0.0
Command-R 35 19.2 3.8 0.0 3.8 3.8 0.0 0.0 3.8 4.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 3.8 0.0 0.0 4.0 0.0 0.0 0.0 0.0
Alfred 40 0.0 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 7.7 3.6 0.0 4.0 0.0 3.8 0.0 0.0
Mixtral 56 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0
Llama 2 70 0.0 3.8 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0
Llama 3 70 38.5 23.1 0.0 7.7 30.8 7.7 0.0 7.7 24.0 11.5 0.0 3.8 14.8 12.1 0.0 11.5 26.9 14.3 0.0 0.0 0.0 7.7 0.0 0.0
Llama 3.1 70 46.2 19.2 0.0 19.2 19.2 19.2 0.0 11.5 36.0 19.2 0.0 3.8 18.5 15.2 0.0 7.7 34.6 10.7 0.0 0.0 4.0 3.8 0.0 0.0
DeepSeek-R1 70 34.6 26.9 0.0 7.7 26.9 26.9 0.0 0.0 24.0 15.4 0.0 0.0 11.1 9.1 0.0 7.7 19.2 14.3 0.0 0.0 8.0 19.2 3.8 0.0
Qwen 2 72 34.6 15.4 3.8 23.1 19.2 7.7 0.0 11.5 8.0 3.8 0.0 0.0 3.7 9.1 0.0 3.8 7.7 0.0 0.0 0.0 0.0 7.7 0.0 3.8
Command-R+ 104 30.8 3.8 3.8 15.4 26.9 3.8 0.0 7.7 4.0 3.8 0.0 3.8 3.7 3.0 0.0 0.0 19.2 3.6 0.0 0.0 4.0 0.0 0.0 0.0
Mistral-Large 123 38.5 19.2 0.0 7.7 15.4 3.8 0.0 3.8 20.0 7.7 0.0 0.0 18.5 0.0 3.8 0.0 26.9 3.6 0.0 0.0 8.0 7.7 0.0 0.0
Mixtral 176 7.7 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 3.6 0.0 0.0 0.0 7.7 0.0 0.0
WizardLM 2 176 19.2 11.5 0.0 7.7 7.7 3.8 0.0 3.8 12.0 3.8 0.0 0.0 3.7 0.0 0.0 0.0 3.8 3.6 0.0 0.0 0.0 7.7 0.0 0.0
Large Models $(>200B)$
Llama 3.1 405 92.3 57.7 80.8 84.6 26.9 38.5 34.6 57.7 76.0 46.2 26.9 50.0 70.4 45.5 65.4 61.5 76.9 39.3 57.7 52.0 44.0 34.6 61.5 50.0
DeepSeek-R1 671 96.2 84.6 96.2 92.3 100 65.4 76.9 92.3 92.0 76.9 42.3 88.5 92.6 57.6 84.6 76.9 80.8 64.3 73.1 72.0 100 100 73.1 57.7
ChatGPT 3.5 unk 100 76.9 100 92.3 100 84.6 69.2 88.5 100 92.3 38.5 96.2 92.6 75.8 80.8 88.5 88.5 60.7 76.9 84.0 76.0 42.3 76.9 65.4
Claude 3.5 S unk 100 80.8 100 88.5 96.2 88.5 88.5 96.2 100 92.3 38.5 96.2 88.9 84.8 84.6 96.2 84.6 71.4 80.8 72.0 64.0 73.1 69.2 69.2
ChatGPT 4o unk 96.2 84.6 100 88.5 100 73.1 80.8 88.5 100 84.6 38.5 96.2 96.3 78.8 80.8 84.6 80.8 57.1 73.1 88.0 64.0 50.0 65.4 61.5
ChatGPT 4o m unk 96.2 73.1 100 92.3 100 84.6 80.8 96.2 72.0 46.2 42.3 80.8 77.8 54.5 50.0 57.7 73.1 39.3 57.7 52.0 48.0 34.6 42.3 50.0
Gemini 2.0 F unk 88.5 76.9 96.2 96.2 96.2 65.4 80.8 96.2 92.0 84.6 42.3 84.6 92.6 66.7 80.8 88.5 84.6 60.7 80.8 76.0 44.0 50.0 61.5 61.5
LeChat unk 80.8 26.9 3.8 57.7 73.1 50.0 38.5 53.8 48.0 26.9 34.6 50.0 74.1 33.3 53.8 53.8 53.8 42.9 26.9 44.0 36.0 26.9 34.6 30.8

### v4

@@ B Full List of LLMs and Reasons for Inclusion
Gemma 3 1 Officially supports only English but manages to generate Luxembourgish in experiments
@@
Gemma 3 4 Supports 140+ languages/ not explicitly supporting Luxembourgish
@@
Gemma 3 12 Supports 140+ languages/ not explicitly supporting Luxembourgish
@@
Gemma 3 27 Supports 140+ languages/ not explicitly supporting Luxembourgish
QWQ 32 Supports 29 languages/ not explicitly supporting Luxembourgish
@@ C Language Exam Prompt
"I will give you a language test for
@@
by [BLANK] and a list of possible
ANSWERS where every possible option is
separated with a comma. You MUST choose
the option out of the provided list that
best fits as replacement for the given
BLANK. As output, ONLY write the chosen
option, nothing else. DO NOT WRITE ANY
ADDITIONAL TEXT."
@@ D Evaluation Metric Prompts for LuxGen
    written without any grammatical or
    spelling mistakes. The more mistakes there
    are, the lower the score should be. A 0
    means that the headline is full of mistakes
    and made up words. Also return 0 if the
    description is not written in
    Luxembourgish at all.
@@ F Fine-grained Exam Results
G Fine-grained Category Description
H Fine-grained Category Annotation Prompt
We use the following prompt to categorise the exam questions:
I have a language test for Luxembourgish
where I would like you to categorise every
question into different types of skill
they are testing. For every question and
every category, determine and mark
whether or not a given question belongs
to a given category, as follows:
0, the question does not belong
to the category
1, the question belongs to the category
Write the output in tabular form like
TSV.
The file I am uploading are some
questions that I already manually
annotated.
Use it as a guideline.
The list of categories and explanations
is as follows:
vocabulary: any question relating
to vocabulary
grammar: any question relating
to grammar
vocabulary_idiom: a question relating
to idioms or metaphors
vocabulary_noun: a vocabulary question
relating to a noun
vocabulary_verb: a vocabulary question
relating to a verb
vocabulary_adjective: a vocabulary
question relating to an adjective
vocabulary_native_word: a vocabulary
question where the correct answer is a
word exclusive to Luxembourgish
grammar_gender: a question relating to determining the gender
of a determinant, adjective, or pronoun
grammar_pronouns: a question relating to pronouns
grammar_verbs:  a question relating to conjugations of verbs
grammar_adjectives:  a question relating to adjectives
grammar_tense: a question relating to the correct use of tenses
grammar_tense_past: a question relating to the past tense
grammar_tense_present: a question relating to the present tense
grammar_verbs_modal_verbs: a question relating to modal verbs
of prepositions or conjunctions
reading_comprehension_true_false: a reading comprehension question where the answer
is either true or false
reading_comprehension_simple: a reading comprehension question that does not require to
understand the entire text excerpt to answer correctly
reading_comprehension_complex: a reading comprehension question that requires a deeper
understanding of the entire text excerpt to answer correctly
I LuxGen Evaluation Results using LLM-as-a-Judge metrics
A1 A2 B1 B2 C1 C2
Model Size Voc. Gra. RC $C$ Voc. Gra. RC $C$ Voc. Gra. RC $C$ Voc. Gra. RC $C$ Voc. Gra. RC $C$ Voc. Gra. RC $C$
Qwen 2 0.5 38.5 26.9 23.1 19.2 34.6 11.5 19.2 23.1 32.0 15.4 23.1 23.1 25.0 17.6 19.2 30.8 16.7 20.7 30.8 20.0 7.7 26.1 11.5 19.2
Llama 3.2 1 15.4 34.6 34.6 19.2 26.9 11.5 19.2 19.2 32.0 46.2 19.2 15.4 21.4 20.6 19.2 23.1 16.7 20.7 11.5 28.0 15.4 8.7 19.2 7.7
Gemma 3 1 38.5 46.2 50.0 46.2 19.2 34.6 50.0 42.3 44.0 42.3 34.6 34.6 14.3 44.1 50.0 53.8 29.2 20.7 26.9 24.0 19.2 39.1 38.5 38.5
Qwen 2 1.5 34.6 38.5 42.3 15.4 38.5 26.9 23.1 23.1 28.0 30.8 19.2 23.1 28.6 38.2 23.1 34.6 25.0 13.8 19.2 28.0 23.1 21.7 23.1 26.9
StableLM 2 1.6 30.8 34.6 7.7 15.4 19.2 23.1 0.0 15.4 20.0 30.8 11.5 23.1 25.0 23.5 11.5 3.8 20.8 24.1 3.8 12.0 23.1 34.8 0.0 11.5
EuroLLM 1.7 19.2 15.4 11.5 23.1 23.1 23.1 15.4 23.1 24.0 11.5 11.5 3.8 21.4 29.4 23.1 19.2 33.3 27.6 15.4 28.0 23.1 13.0 23.1 11.5
Gemma 2 2 38.5 46.2 38.5 34.6 38.5 23.1 34.6 34.6 44.0 19.2 30.8 30.8 17.9 32.4 30.8 50.0 29.2 31.0 19.2 32.0 38.5 26.1 23.1 26.9
Llama 3.2 3 34.6 34.6 46.2 34.6 38.5 23.1 26.9 53.8 32.0 26.9 23.1 23.1 32.1 26.5 42.3 42.3 37.5 24.1 23.1 40.0 15.4 34.8 30.8 19.2
Phi 3 3.8 26.9 46.2 57.7 50.0 34.6 23.1 38.5 50.0 24.0 23.1 23.1 38.5 25.0 26.5 15.4 38.5 33.3 24.1 23.1 32.0 23.1 8.7 19.2 19.2
Phi 3.5 3.8 34.6 50.0 30.8 42.3 26.9 30.8 15.4 38.5 8.0 19.2 11.5 7.7 14.3 8.8 11.5 7.7 12.5 6.9 7.7 4.0 7.7 4.3 15.4 15.4
Gemma 3 4 65.4 26.9 69.2 57.7 38.5 50.0 34.6 50.0 48.0 34.6 30.8 46.2 57.1 29.4 42.3 61.5 33.3 17.2 53.8 44.0 34.6 39.1 30.8 23.1
Qwen 2 7 42.3 26.9 69.2 53.8 23.1 38.5 38.5 46.2 28.0 23.1 15.4 61.5 28.6 35.3 38.5 42.3 58.3 24.1 38.5 40.0 34.6 39.1 15.4 38.5
Llama 2 7 26.9 34.6 53.8 30.8 46.2 23.1 34.6 50.0 24.0 26.9 26.9 19.2 46.4 32.4 38.5 42.3 29.2 17.2 19.2 36.0 11.5 26.1 38.5 30.8
WizardLM 2 7 42.3 34.6 65.4 34.6 50.0 34.6 26.9 42.3 24.0 15.4 15.4 30.8 42.9 23.5 30.8 30.8 29.2 37.9 23.1 24.0 26.9 47.8 11.5 34.6
Mistral 7 30.8 26.9 46.2 38.5 23.1 30.8 42.3 46.2 40.0 34.6 26.9 34.6 28.6 23.5 26.9 46.2 41.7 27.6 34.6 36.0 23.1 26.1 30.8 26.9
Aya-23 8 38.5 38.5 46.2 38.5 30.8 19.2 30.8 42.3 36.0 19.2 26.9 38.5 25.0 32.4 26.9 42.3 29.2 41.4 23.1 32.0 30.8 26.1 38.5 30.8
Llama 3.1 8 50.0 61.5 57.7 57.7 30.8 30.8 26.9 46.2 16.0 38.5 19.2 57.7 32.1 23.5 53.8 26.9 20.8 20.7 50.0 32.0 30.8 26.1 30.8 46.2
Llama 3 8 42.3 42.3 50.0 46.2 30.8 34.6 26.9 50.0 40.0 30.8 30.8 38.5 25.0 38.2 34.6 46.2 45.8 17.2 26.9 28.0 23.1 39.1 42.3 23.1
DeepSeek-R1 8 69.2 42.3 57.7 46.2 46.2 38.5 26.9 46.2 48.0 42.3 30.8 34.6 35.7 38.2 46.2 38.5 29.2 24.1 50.0 36.0 15.4 21.7 30.8 46.2
GLM 4 9 57.7 42.3 69.2 50.0 34.6 19.2 42.3 69.2 56.0 15.4 26.9 53.8 60.7 26.5 50.0 50.0 54.2 34.5 46.2 44.0 38.5 26.1 26.9 50.0
EuroLLM 9 53.8 30.8 57.7 34.6 50.0 38.5 42.3 50.0 32.0 42.3 30.8 26.9 39.3 23.5 34.6 42.3 41.7 31.0 23.1 32.0 23.1 26.1 26.9 26.9
Gemma 2 9 76.9 50.0 61.5 46.2 61.5 38.5 30.8 57.7 48.0 26.9 34.6 46.2 64.3 38.2 53.8 42.3 58.3 20.7 46.2 40.0 34.6 30.4 42.3 30.8
Mistral-Nemo 12 61.5 50.0 53.8 46.2 34.6 26.9 38.5 46.2 36.0 15.4 23.1 38.5 28.6 23.5 38.5 46.2 37.5 24.1 26.9 32.0 19.2 26.1 26.9 11.5
StableLM 2 12 42.3 15.4 30.8 23.1 4.62 19.2 34.6 42.3 44.0 11.5 19.2 30.8 21.4 17.6 50.0 30.8 37.5 17.2 19.2 40.0 19.2 21.7 30.8 23.1
Gemma 3 12 76.9 30.8 73.1 69.2 80.8 42.3 46.2 65.4 64.0 46.2 42.3 61.5 64.3 29.4 57.7 57.7 54.2 24.1 61.5 52.0 38.5 21.7 46.2 53.8
Llama 2 13 34.6 38.5 15.4 23.1 19.2 11.5 15.4 11.5 8.0 34.6 19.2 30.8 32.1 20.6 30.8 23.1 25.0 24.1 23.1 20.0 34.6 30.4 23.1 19.2
Phi 3 14 42.3 42.3 50.0 46.2 38.5 42.3 30.8 42.3 40.0 38.5 30.8 46.2 46.4 35.3 34.6 42.3 25.0 34.5 30.8 20.0 11.5 21.7 19.2 34.6
Phi 4 14 50.0 50.0 65.4 50.0 65.4 42.3 50.0 50.0 36.0 42.3 34.6 57.7 57.1 44.1 46.2 30.8 41.7 24.1 34.6 40.0 26.9 17.4 42.3 46.2
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 46.2 34.6 69.2 57.7 57.7 19.2 34.6 73.1 44.0 19.2 15.4 46.2 42.9 38.2 30.8 46.2 45.8 31.0 38.5 40.0 11.5 34.8 19.2 38.5
Gemma 2 27 65.4 46.2 80.8 73.1 73.1 50.0 57.7 73.1 68.0 38.5 42.3 69.2 60.7 26.5 57.7 42.3 62.5 24.1 50.0 68.0 30.8 34.8 53.8 46.2
Gemma 3 27 76.9 46.2 84.6 76.9 96.2 57.7 61.5 73.1 76.0 61.5 42.3 69.2 67.9 52.9 65.4 57.7 75.0 48.3 61.5 60.0 34.6 43.5 50.0 50.0
QWQ 32 80.8 46.2 61.5 53.8 61.5 42.3 30.8 61.5 64.0 42.3 34.6 65.4 50.0 38.2 53.8 50.0 62.5 20.7 34.6 40.0 38.5 39.1 34.6 65.4
Aya-23 35 34.6 34.6 61.5 42.3 46.2 15.4 38.5 50.0 40.0 46.2 30.8 42.3 57.1 17.6 46.2 26.9 25.0 31.0 34.6 44.0 34.6 39.1 30.8 34.6
Command-R 35 42.3 53.8 61.5 46.2 46.2 19.2 38.5 50.0 36.0 26.9 30.8 38.5 35.7 35.3 50.0 50.0 37.5 27.6 38.5 40.0 38.5 43.5 46.2 23.1
Alfred 40 30.8 38.5 61.5 38.5 50.0 30.8 26.9 50.0 24.0 30.8 26.9 34.6 28.6 32.4 30.8 26.9 45.8 31.0 34.6 36.0 26.9 30.4 26.9 34.6
Mixtral 56 50.0 38.5 61.5 50.0 50.0 19.2 42.3 53.8 48.0 19.2 26.9 42.3 50.0 38.2 34.6 34.6 50.0 13.8 23.1 44.0 34.6 34.8 30.8 19.2
Llama 2 70 23.1 34.6 57.7 38.5 34.6 26.9 30.8 42.3 24.0 26.9 26.9 34.6 39.3 35.3 30.8 53.8 25.0 24.1 26.9 28.0 26.9 21.7 46.2 61.5
Llama 3 70 61.5 34.6 65.4 50.0 73.1 30.8 38.5 53.8 36.0 38.5 38.5 42.3 42.9 26.5 38.5 34.6 58.3 31.0 34.6 32.0 46.2 52.2 26.9 50.0
Llama 3.1 70 57.7 38.5 69.2 57.7 84.6 34.6 42.3 65.4 56.0 53.8 34.6 53.8 71.4 44.1 46.2 50.0 58.3 48.3 42.3 40.0 34.6 56.5 38.5 38.5
DeepSeek-R1 70 76.9 53.8 65.4 61.5 69.2 30.8 30.8 73.1 68.0 61.5 42.3 53.8 67.9 35.3 50.0 69.2 62.5 41.4 38.5 56.0 42.3 30.4 42.3 50.0
Qwen 2 72 65.4 53.8 65.4 57.7 50.0 34.6 42.3 57.7 44.0 34.6 38.5 46.2 57.1 41.2 57.7 53.8 62.5 24.1 42.3 40.0 30.8 30.4 46.2 53.8
Command-R+ 104 57.7 38.5 61.5 53.8 50.0 19.2 42.3 42.3 52.0 30.8 30.8 46.2 42.9 47.1 42.3 46.2 45.8 20.7 34.6 28.0 30.8 30.4 30.8 38.5
Mistral-Large 123 84.6 57.7 73.1 50.0 65.4 34.6 38.5 61.5 44.0 42.3 30.8 53.8 75.0 38.2 50.0 50.0 66.7 34.5 26.9 48.0 53.8 39.1 34.6 50.0
Mixtral 176 57.7 57.7 73.1 61.5 57.7 50.0 38.5 50.0 44.0 30.8 38.5 53.8 60.7 35.3 34.6 38.5 41.7 34.5 38.5 44.0 34.6 34.8 30.8 57.7
WizardLM 2 176 53.8 50.0 65.4 65.4 42.3 30.8 30.8 57.7 32.0 38.5 38.5 53.8 50.0 41.2 46.2 34.6 37.5 31.0 30.8 44.0 26.9 34.8 38.5 38.5
Large Models $(>200B)$
Llama 3.1 405 92.3 57.7 80.8 84.6 26.9 38.5 34.6 57.7 72.0 46.2 26.9 50.0 67.9 47.1 65.4 61.5 79.2 37.9 57.7 52.0 42.3 30.4 61.5 50.0
DeepSeek-R1 671 96.2 84.6 96.2 92.3 100.0 65.4 76.9 92.3 88.0 76.9 42.3 88.5 89.3 58.8 84.6 76.9 83.3 65.5 73.1 72.0 96.2 100.0 73.1 57.7
ChatGPT 3.5 unk 100.0 76.9 100.0 92.3 100.0 84.6 69.2 88.5 96.0 92.3 38.5 96.2 89.3 76.5 80.8 88.5 87.5 58.6 76.9 84.0 76.9 43.5 76.9 65.4
Claude 3.5 S unk 100.0 80.8 100.0 88.5 96.2 88.5 88.5 96.2 96.0 92.3 38.5 96.2 85.7 82.4 84.6 96.2 87.5 72.4 80.8 72.0 61.5 78.3 69.2 69.2
ChatGPT 4o unk 96.2 84.6 100.0 88.5 100.0 73.1 80.8 88.5 96.0 84.6 38.5 96.2 92.9 79.4 80.8 84.6 83.3 55.2 73.1 88.0 61.5 52.2 65.4 61.5
ChatGPT 4o m unk 96.2 73.1 100.0 92.3 100.0 84.6 80.8 96.2 68.0 46.2 42.3 80.8 75.0 55.9 50.0 57.7 75.0 37.9 57.7 52.0 46.2 34.8 42.3 50.0
Gemini 2.0 F unk 88.5 76.9 96.2 96.2 96.2 65.4 80.8 96.2 88.0 84.6 42.3 84.6 89.3 67.6 80.8 88.5 87.5 62.1 80.8 76.0 42.3 52.2 61.5 61.5
LeChat unk 80.8 26.9 3.8 57.7 73.1 50.0 38.5 53.8 44.0 26.9 34.6 50.0 71.4 32.4 53.8 53.8 54.2 41.4 26.9 44.0 34.6 30.4 34.6 30.8